OpenStudy (anonymous):

What is the exact value of √50/3√15

OpenStudy (anonymous):

9.128709292

hero (hero):

Hi Sarah

OpenStudy (anonymous):

\[5\sqrt{2}/3\sqrt{15}\]

OpenStudy (mimi_x3):

\[\large \frac{\sqrt{50}}{3\sqrt{15}} \]?

hero (hero):

√50 ÷ 3√15

hero (hero):

rationalizing the denominator would be a good first step

OpenStudy (earthcitizen):

\[(5\sqrt{2}/3\sqrt{15})*(\sqrt{15}/\sqrt{15})\]

OpenStudy (earthcitizen):

\[\sqrt{30}/9\]